### v1

# arXiv:2603.01390v1[math.RT] 02 Mar 2026

arXiv:2603.01390v1[math.RT] 02 Mar 2026
Duflo-Serganova functors and Brundan-Goodwin’s parabolic inductions
Shunsuke Hirota
Duflo–Serganova functors play a central role in the representation theory of Lie superalgebras and have been studied extensively, especially in the finite-dimensional setting. On the other hand, in the general linear case, Brundan–Goodwin observed that certain distinguished parabolic induction functors are particularly compatible with the principal Whittaker coinvariants functor $H_{0}$, and these functors are crucial in their theory.
In this paper we show that, for certain rank-one DS functors, the image under DS can be computed explicitly for a class of infinite-dimensional modules which can be realized as images of these parabolic induction functors. In particular, for $\mathfrak{b}$-Verma supermodules attached to a suitable class of Borel subalgebras $\mathfrak{b}$ and a family of modules whose $H_{0}$-images are tensor products of one-dimensional evaluation modules for the super Yangian, we obtain an elegant explicit formula.
1 Introduction
Highest weight representations, together with BGG category $\mathcal{O}$, are among the most fundamental and central objects in the representation theory of semisimple Lie algebras. In this area, endofunctors attached to roots—such as translation functors—have played a crucial role for decades [21]. Representation theory of basic classical Lie superalgebras, in particular for the general linear Lie superalgebra $\mathfrak{gl}(m|n)$ [6], provides a natural extension of the classical theory and exhibits phenomena absent in the Lie algebra case, such as the fact that Borel subalgebras need not be conjugate. Nevertheless, many of the classical constructions still make sense. For instance, twisting functors associated with even roots have been studied extensively [10]. It is therefore natural to ask whether there exist similarly well-behaved functors associated with odd roots, a question which is expected to reveal phenomena specific to the super setting.
The Duflo–Serganova functor [12] is widely regarded as the canonical example—and perhaps the only currently known example—of a “good” functor attached to an odd root. For $\mathfrak{gl}(m|n)$ it maps representations to those of a smaller Lie superalgebra $\mathfrak{gl}(m-r|n-r)$, thereby reducing complexity. This functor is symmetric monoidal and has left a deep imprint on the theory of finite-dimensional representations; we do not attempt to survey this vast literature here, and instead refer the reader to the survey [14]. For the purposes of this paper, we single out one striking instance: Heidersdorf–Weissauer [16] determine the images of finite-dimensional irreducible $\mathfrak{gl}(n|n)$-modules under Duflo–Serganova functors in terms of Khovanov arc diagrams.
In contrast, much less is known on the infinite-dimensional side. In [11], Coulembier and Serganova study which Verma modules attached to the distinguished Borel are annihilated by Duflo–Serganova functors, and apply this to homological-algebraic questions (these results have been generalized to arbitral Borel subalgebras in part by the author in [19]). Hoyt–Penkov–Serganova [20] show that Duflo–Serganova functors preserve category $\mathcal{O}$ and formulate further conjectures on their behavior. Taken together, these developments suggest that it is natural to investigate the behavior of Duflo–Serganova functors on category $\mathcal{O}$ beyond the finite-dimensional setting. Nevertheless, to the best of the author’s knowledge, concrete nontrivial explicit computations for infinite-dimensional highest weight modules have not appeared. The aim of this paper is to exhibit a class of infinite-dimensional highest weight representations of $\mathfrak{gl}(n|n)$ for which explicit computations with a rank-one Duflo–Serganova functor become possible.
Hypercubic decomposition of Verma modules. A first candidate would be the $\mathfrak{b}$-Verma modules for a Borel $\mathfrak{b}$. However, in view of the fact that DS functors are middle exact, it is natural to decompose $\mathfrak{b}$-Verma modules along the directions of odd $\mathfrak{b}$-simple roots. In [19], the author formalized this idea by realizing, inside $\mathcal{O}_{\mathfrak{gl}(n|n)}$, certain abelian subcategories which are equivalent to blocks of $\mathcal{O}_{\mathfrak{gl}(1|1)^{\oplus n}}$ and which contain the relevant modules. Variants of this idea already appear in the literature, for instance in work of Cheng–Lam–Wang [8] and Serganova [24]. Pursuing this direction leads to a remarkable family of hypercube Borels, for which some of the associated $\mathfrak{b}$-Verma modules can be realized inside an abelian subcategory equivalent to the maximal atypical block of $\mathcal{O}_{\mathfrak{gl}(1|1)^{\oplus n}}$. In this sense, this situation provides the finest possible decomposition of $\mathfrak{b}$-Verma modules.
For example, already for $\mathfrak{gl}(3|3)$ one can visualize the collection of Borel subalgebras with fixed even part as a finite graph whose vertices are Borels and whose edges correspond to odd reflections. In this picture there is a distinguished “cube” in the middle, and the remarkable family of hypecube Borels refers to the Borel subalgebras corresponding to the vertices of this central cube. Note that here the distinguished Borel corresponds to $()$.
[…]
Such Borels have appeared in several places; for example, they are used essentially in Cheng–Lam’s proof of super duality [9].
Parabolic induction and Whittaker coinvariant functor.
Studying the decomposition of Verma modules for hypercube Borels is equivalent to considering the parabolic induction functor of Brundan–Goodwin [4], which sends $\mathfrak{gl}(1|1)^{\oplus n}$-modules to $\mathfrak{gl}(n|n)$-modules. The relevance of these parabolic inductions is explained by the principal Whittaker coinvariant functor $H_{0}$, which sends objects of category $\mathcal{O}$ to modules over the (nontrivial!) principal $W$-superalgebra. Generalizing [5], Brown–Brundan–Goodwin [3] show that the principal $W$-superalgebra can be realized as a quotient of the super Yangian $Y(\mathfrak{gl}(1|1))$.
Brundan–Goodwin study $H_{0}$ by exploiting the fact that the above parabolic induction images are sent, under $H_{0}$, to tesor product of evaluation modules for the super Yangian. In particular, they show that Verma modules for the distinguished Borel are mapped to Verma modules on the $W$-superalgebra side, defined via the Drinfeld presentation of super Yangians, and that the Whittaker coinvariant functor $H_{0}$ provides a realization of the Soergel functor. Thus these parabolic induction functors, and their images, play an essential role in the study of category $\mathcal{O}$; we return to this point in §3.
What we prove. In §5 we show that, for a certain rank-one Duflo–Serganova functor, the image under DS can be computed explicitly on a class of modules containing the images of the parabolic induction functors considered by Brundan–Goodwin. In particular, in a setting which includes the case where $\mathfrak{b}$ is a hypercube Borel and $e_{\alpha}=e_{1,n+1}$, we prove the following conjecture.
Conjecture 1.1.
Let $\mathfrak{g}=\mathfrak{gl}(n|n)$, let $\mathfrak{b}$ be a Borel subalgebra, and let $\alpha$ be a $\mathfrak{b}$-simple odd root. Then for $\lambda\in\Lambda$ one has
$\operatorname{DS}_{e_{\alpha}}\bigl(M^{\mathfrak{b}}(\lambda)\bigr)\ \cong\ \begin{cases}M^{\mathfrak{b}_{e_{\alpha}}}\bigl(\operatorname{pr}_{\alpha}(\lambda)\bigr)\ \oplus\ \Pi\,M^{\mathfrak{b}_{e_{\alpha}}}\bigl(\operatorname{pr}_{\alpha}(\lambda)\bigr),&\text{if }(\lambda,\alpha)=0,\\[5.69054pt] 0,&\text{if }(\lambda,\alpha)\neq 0.\end{cases}$
Here $\operatorname{pr}_{\alpha}$ is the natural projection on weight lattices associated with $\alpha$, defined in §6.
In §6 we verify this conjecture in the case of $\mathfrak{gl}(2|2)$ by direct calculations using Hinich’s lemma.
As suggested by the conjecture, our computations for $\mathfrak{b}$-Verma modules are not expected to depend essentially on our particular choice of Borel subalgebras. However, an important feature of our hypercube Borels is the following result.
Theorem 1.2 (Theorem 5.9).
Let $\mathfrak{g}=\mathfrak{gl}(n|n)$ and let $\operatorname{BG}_{n|n}(\lambda)$ be the module obtained by applying the Brundan–Goodwin parabolic induction functor to an irreducible $\mathfrak{gl}(1|1)^{\oplus n}$-module with highest weight $\lambda$. If $H_{0}\operatorname{BG}_{n|n}(\lambda)$ is one-dimensional, then, up to parity,
$\operatorname{DS}_{e_{1,n+1}}\operatorname{BG}_{n|n}(\lambda)\ \cong\ \operatorname{BG}_{n-1|n-1}\bigl(\operatorname{pr}_{\alpha}(\lambda)\bigr),$
and moreover $H_{0}\operatorname{BG}_{n-1|n-1}\bigl(\operatorname{pr}_{\alpha}(\lambda)\bigr)$ is also one-dimensional.
Remarks. Although we expect that our results can be generalized to the $\mathfrak{gl}(m|n)$ setting, we believe that, for the development of this area, it is important to focus on $\mathfrak{gl}(n|n)$. Working with $\mathfrak{gl}(n|n)$ brings a certain simplicity and beauty. For instance, in the $\mathfrak{gl}(n|n)$ case one does not need to consider shifts of the super Yangian [3]. Also, a number of existing works suggest that phenomena for $\mathfrak{gl}(m|n)$ often reduce to, or are controlled by, the $\mathfrak{gl}(n|n)$ case; see for example [16, 17, 25].
The main objects of study in this paper, namely the modules $\operatorname{BG}(\lambda)$, can be realized as images of homomorphisms between Verma modules attached to Borel subalgebras lying at opposite vertices of the central hypercube. On the other hand, in [18] the images of homomorphisms between Verma modules attached to the distinguished and anti-distinguished Borels were studied, and were used in a certain generalization of a classical BGG-type realization. In a sense, these two constructions may be viewed as lying at opposite extremes.
Finally, we note that, from the perspective of Nichols algebras of diagonal type and their classification via Weyl groupoids (a.k.a. generalized root systems), changes of Borel are of independent interest (see [2, 1, 15]. We believe that the results of this paper (Theorems 3.13 and 5.9 provide further evidence for the significance of such changes of Borel.
1.1 Acknowledgments
This work was supported by the Japan Society for the Promotion of Science (JSPS) through the Research Fellowship for Young Scientists (DC1), Grant Number JP25KJ1664.
2 basic facts about $\mathfrak{gl}(n|n)$
In this section, we summarize basic facts about $\mathfrak{gl}(n|n)$ and its representation theory. All material in this section is well known.
Let the base field $k$ be an algebraically closed field of characteristic 0. Let $\mathbf{Vec}$ denote the category of vector spaces, and $\mathbf{sVec}$ the category of supervector spaces. Let $\Pi:\mathbf{sVec}\to\mathbf{sVec}$ be the parity–shift functor. Let $F:\mathbf{sVec}\to\mathbf{Vec}$ be the monoidal functor that forgets the $\mathbb{Z}/2\mathbb{Z}$–grading. Note that $F$ is not a symmetric monoidal functor. In what follows, whenever we refer to the dimension of a supervector space, we mean this total (ordinary) dimension:
$\dim V:=\dim_{k}F(V).$
From now on, we will denote by $\mathfrak{g}$ a finite dimensional Lie superalgebra. We consider the category $\mathfrak{g}\text{-sMod}$, where morphisms respects $\mathbb{Z}/2\mathbb{Z}$-grading. (This is the module category of a monoid object in $\mathfrak{g}\text{-sMod}$ in the sense of [13].)
Let $V=V_{\overline{0}}\oplus V_{\overline{1}}$ be a $\mathbb{Z}/2\mathbb{Z}$-graded vector space, where $V_{\overline{0}}$ (the even part) is spanned by $v_{1},\dots,v_{n}$ and $V_{\overline{1}}$ (the odd part) is spanned by $v_{n+1},\dots,v_{2n}$.
The space $\operatorname{End}(V)$ is spanned by basis elements $e_{ij}$, defined by: $e_{ij}\cdot v_{k}=\delta_{jk}v_{i}.$
The general linear Lie superalgebra $\mathfrak{gl}(n|n)$ is defined as the Lie superalgebra spanned by all $e_{ij}$ with $1\leq i,j\leq 2n$, under the supercommutator: $e_{ij},e_{kl}]=e_{ij}e_{kl}-(-1)^{|e_{ij}||e_{kl}|}e_{kl}e_{ij},$) where $|e_{ij}|=\overline{0}$ if $e_{ij}$ acts within $V_{\overline{0}}$ or $V_{\overline{1}}$ (even), and $|e_{ij}|=\overline{1}$ if it maps between $V_{\overline{0}}$ and $V_{\overline{1}}$ (odd).
The even part is given by $\mathfrak{g}_{\overline{0}}=\mathfrak{gl}(n)\oplus\mathfrak{gl}(n).$
We fix the standard Cartan subalgebra $\mathfrak{h}:=\bigoplus_{1\leq i\leq 2n}ke_{ii}.$ Define linear functionals $\varepsilon_{1},\dots,\varepsilon_{2n}\in\mathfrak{h}^{*}$ by requiring that $\varepsilon_{i}(e_{jj})=\delta_{ij}$ for $1\leq i,j\leq 2n$. Define $\delta_{i}=\varepsilon_{n+i}$ for $1\leq i\leq n$. The non-degenerate symmetric bilinear form (,)(\,,\,) on is defined as follows: $(\varepsilon_{i},\varepsilon_{j})=(-1)^{|v_{i}|}\delta_{i,j}$
The root space $\mathfrak{g}_{\alpha}$ associated with $\alpha\in\mathfrak{h}^{*}$ is defined as $\mathfrak{g}_{\alpha}:=\{\,x\in\mathfrak{g}\mid[h,x]=\alpha(h)x,\text{ for all }h\in\mathfrak{h}\,\}.$ Then we have $\mathfrak{g}_{\varepsilon_{i}-\varepsilon_{j}}=ke_{ij}.$
The set of roots $\Delta$ is defined as $\Delta:=\{\,\alpha\in\mathfrak{h}^{*}\mid\mathfrak{g}_{\alpha}\neq 0\,\}\setminus\{0\}.$
We have a root space decomposition of $\mathfrak{g}$ with respect to $\mathfrak{h}$:
$\mathfrak{g}=\mathfrak{h}\oplus\bigoplus_{\alpha\in\Delta}\mathfrak{g}_{\alpha},\qquad\text{and}\quad\mathfrak{g}_{0}=\mathfrak{h}.$
we define The sets of all even roots and odd roots as follows: $\Delta_{\overline{0}}=\{\,\varepsilon_{i}-\varepsilon_{j},\ \delta_{i}-\delta_{j}\mid i\neq j\,\},$ $\Delta_{\overline{1}}=\{\,\varepsilon_{i}-\delta_{j}\mid 1\leq i\leq n,\ 1\leq j\leq n\,\}.$ Note that for any odd root $\alpha\in\Delta_{\bar{1}}$ we have $(\alpha,\alpha)=0$; i.e., every odd root is isotropic .
Definition 2.1.
A subset $\Delta^{+}\subset\Delta$ is called a positive system if it satisfies:
  1. 1.
    $\Delta=\Delta^{+}\cup(-\Delta^{+})$ (disjoint union),
  2. 2.
    For any $\alpha,\beta\in\Delta^{+}$, if $\alpha+\beta\in\Delta$, then $\alpha+\beta\in\Delta^{+}$.
Let $\Delta^{+}\subset\Delta$ be a positive system. The corresponding Borel subalgebra $\mathfrak{b}\subset\mathfrak{g}$ is defined as $\mathfrak{b}:=\mathfrak{h}\oplus\bigoplus_{\alpha\in\Delta^{+}}\mathfrak{g}_{\alpha},$ where $\mathfrak{g}_{\alpha}\subset\mathfrak{g}$ is the root space corresponding to $\alpha\in\Delta$. Choose natural root vectors $e_{\alpha}\in\mathfrak{g}_{\alpha}$, $e_{-\alpha}\in\mathfrak{g}_{\-\alpha}$ Then for any $\lambda\in\mathfrak{h}^{*}$, $\lambda([e_{\alpha},e_{-\alpha}])=(\lambda,\alpha).$
The sets of positive roots, even positive roots, and odd positive roots corresponding to $\mathfrak{b}$ are denoted by $\Delta^{\mathfrak{b},+}$, $\Delta_{\bar{0}}^{\mathfrak{b},+}$, and $\Delta_{\bar{1}}^{\mathfrak{b},+}$, respectively.
The Weyl group $S_{n}\times S_{n}$ of $\mathfrak{g}_{\bar{0}}=\mathfrak{gl}(n)\oplus\mathfrak{gl}(n)$ acts on $\mathfrak{h}^{*}$ by permuting the $\varepsilon_{i}$’s and the $\delta_{j}$’s separately. In particular, the even positive root systems are all conjugate under $S_{n}\times S_{n}$. We therefore fix the standard even Borel $\mathfrak{b}^{\mathrm{st}}_{\bar{0}}$ (block upper triangular in $\mathfrak{g}_{\bar{0}}$), with even positive root system
$\Delta_{\bar{0}}^{\mathrm{st},+}=\{\ \varepsilon_{i}-\varepsilon_{j}\mid 1\leq i<j\leq n\ \}\ \cup\ \{\ \delta_{p}-\delta_{q}\mid 1\leq p<q\leq n\ \}.$
With this choice fixed, positive systems whose even part equals $\Delta_{\bar{0}}^{\mathrm{st},+}$ are classified by $\varepsilon\delta$–sequences, equivalently by $(n|n)$-shuffle
$\tau\in\mathrm{Sh}(n|n):=\bigl\{\,w\in S_{2n}\ \big|\ w(1)<\cdots<w(n)\ \text{and}\ w(n+1)<\cdots<w(2n)\,\bigr\}.$
Given $\tau$, set
$\Delta^{+}(\tau)=\{\ \varepsilon_{i}-\varepsilon_{j}\mid\tau(i)<\tau(j)\ \}$
yielding a bijection $\tau\leftrightarrow\Delta^{+}(\tau)$; hence there are $\frac{(2n)!}{n!n!}$ positive systems with standard even one.
Definition 2.2.
A partition is a weakly decreasing finite sequence of positive integers; we allow the empty partition (). For example, we write $5+3+3+1$ as $(53^{2}1)$. We identify partitions with Young diagrams in French notation (rows increase downward). A box has coordinates $(i,j)$ with $i$ the column (from the left) and $j$ the row (from the bottom).
Each $\varepsilon\delta$–sequence with $n$ symbols $\varepsilon$ and $n$ symbols $\delta$ determines a lattice path from $(n,0)$ to $(0,n)$ (left step for $\varepsilon$, up step for $\delta$); the region weakly southeast of the path inside $(n^{n})$ is a Young diagram fitting in the $n\times n$ rectangle $(n^{n})$, and this gives a bijection between $\varepsilon\delta$–sequences and the set of Young diagrams fitting in the $n\times n$ rectangle $(n^{n})$.
We fix the standard even Borel subalgebra $\mathfrak{b}^{\mathrm{st}}_{\bar{0}}\subset\mathfrak{g}_{\bar{0}}$ corresponding to the even positive system $\Delta_{\bar{0}}^{\mathrm{st},+}$. We write
$L(n,n):=\Bigl\{\ \mathfrak{b}:\text{Borel subalgebra}\ \Bigm|\ \mathfrak{b}_{\bar{0}}=\mathfrak{b}^{\mathrm{st}}_{\bar{0}}\ \Bigr\}.$
In what follows, we represent a Borel subalgebra with standard even Borel subalgebra by a partition. In particular, we call () the distinguished Borel subalgebra, and $(n^{n})$ the anti-distinguished Borel subalgebra.
Definition 2.3.
A weight $\lambda\in\mathfrak{h}^{*}$ is integral if
$\lambda=\sum_{i=1}^{n}a_{i}\,\varepsilon_{i}+\sum_{j=1}^{n}b_{j}\,\delta_{j}\qquad\text{with }a_{i},b_{j}\in\mathbb{Z}.$
Write the set of integral weights as $\Lambda$.
We denote by $s\mathcal{W}$ the full subcategory of $\mathfrak{g}$-modules which are $\mathfrak{h}$-semisimple with integral weights and finite-dimensional weight spaces. A module $M\in s\mathcal{W}$ admits a weight space decomposition
$M=\bigoplus_{\lambda\in\Lambda}M_{\lambda},\qquad M_{\lambda}:=\{\,v\in M\mid h\cdot v=\lambda(h)v\text{ for all }h\in\mathfrak{h}\,\}.$
Lemma 2.4 (See also [6], Lemma 2.2).
We can choose $\operatorname{par}\in\operatorname{Map}(\Lambda,\mathbb{Z}/2\mathbb{Z})$ such that
$\mathcal{W}:=\{\,M\in s\mathcal{W}\mid\deg M_{\lambda}=\operatorname{par}(\lambda)\text{ for }\lambda\in\Lambda,\ M_{\lambda}\neq 0\,\}$
forms a Serre subcategory, $\mathcal{W}$ contains the trivial module, and $s\mathcal{W}=\mathcal{W}\oplus\Pi\mathcal{W}$.
Similarly, for $\mathfrak{g}_{\bar{0}}$ we introduce the category $s\mathcal{W}_{\bar{0}}$, and define the category $\mathcal{W}_{\bar{0}}$ so as to be compatible with the restriction functor $\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}$.
Definition 2.5.
For a module $M$ in the category $\mathcal{W}$, the character $\operatorname{ch}M$ is the formal sum
$\operatorname{ch}M:=\sum_{\lambda\in\Lambda}\dim M_{\lambda}\,e^{\lambda}.$
Let $L_{\mathfrak{h}}(\lambda)$ be the one-dimensional even $\mathfrak{h}$-module of weight $\lambda\in\Lambda$. We define the $\mathfrak{b}$-Verma module by
$M^{\mathfrak{b}}(\lambda):=\Pi^{\operatorname{par}(\lambda)}\,\operatorname{Ind}_{\mathfrak{b}}^{\mathfrak{g}}\,\operatorname{Infl}_{\mathfrak{h}}^{\mathfrak{b}}L_{\mathfrak{h}}(\lambda).$
Here the parity shift $\Pi^{\operatorname{par}(\lambda)}$ is chosen so that $M^{\mathfrak{b}}(\lambda)\in\mathcal{W}$. Its simple top is denoted by $L^{\mathfrak{b}}(\lambda)$.
Similarly, for the even part $\mathfrak{g}_{\bar{0}}$, the corresponding Verma module and simple module are denoted by $M_{\bar{0}}(\lambda)$ and $L_{\bar{0}}(\lambda)$, respectively.
We define the Berezinian weight
$\operatorname{ber}:=\varepsilon_{1}+\cdots+\varepsilon_{n}-(\delta_{1}+\cdots+\delta_{n}).$
A weight $\lambda$ is orthogonal to all roots if and only if $\lambda=t\,\operatorname{ber}$ for some $t\in k$.
Definition 2.6 (integral Weyl vectors [6]).
Define the vectors
$\rho_{\bar{0}}:=\frac{1}{2}\sum_{\beta\in\Delta_{\bar{0}}^{+}}\beta,\qquad\rho_{\bar{1}}^{\mathfrak{b}}:=\frac{1}{2}\sum_{\gamma\in\Delta_{\bar{1}}^{\mathfrak{b}+}}\gamma,$
and write $\rho_{\bar{1}}:=\rho_{\bar{1}}^{()}$. We have $\rho_{\bar{1}}=-\rho_{\bar{1}}^{(n^{n})}$.
Let the integral Weyl vector be
$\rho:=-\varepsilon_{2}-2\varepsilon_{3}-\cdots-(n-1)\varepsilon_{n}+(n-1)\delta_{1}+(n-2)\delta_{2}+\cdots+\delta_{n-1}.$
One checks that
$\rho=\rho_{\bar{0}}-\rho_{\bar{1}}^{()}-\frac{1}{2}(2n-1)\,\operatorname{ber}.$
For a Borel $\mathfrak{b}$, set
$\rho^{\mathfrak{b}}:=\rho_{\bar{0}}-\rho_{\bar{1}}^{\mathfrak{b}}+\frac{1}{2}\,\operatorname{ber}.$
It is convenient to encode an integral weight $\lambda$ by the $n\mid n$-tuple
$(\lambda_{1},\dots,\lambda_{n}\mid\lambda_{n+1},\dots,\lambda_{2n})\in\mathbb{Z}^{2n},\qquad\lambda_{i}:=(\lambda+\rho,\varepsilon_{i}).$
We write $M^{\mathfrak{b}}(\lambda_{1},\dots,\lambda_{n}\mid\lambda_{n+1},\dots,\lambda_{2n}):=M^{\mathfrak{b}}(\lambda+\rho-\rho^{\mathfrak{b}})$. Similarly, we write $L^{\mathfrak{b}}(\lambda_{1},\dots,\lambda_{n}\mid\lambda_{n+1},\dots,\lambda_{2n}):=L^{\mathfrak{b}}(\lambda+\rho-\rho^{\mathfrak{b}})$ and $P^{\mathfrak{b}}(\lambda_{1},\dots,\lambda_{n}\mid\lambda_{n+1},\dots,\lambda_{2n}):=P^{\mathfrak{b}}(\lambda+\rho-\rho^{\mathfrak{b}})$.
For an integral weight $\lambda=(\lambda_{1},\dots,\lambda_{n}\mid\lambda_{n+1},\dots,\lambda_{2n})$, $\lambda$ (or $L^{()}(\lambda)$) is antidominant iff
$\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{n}\qquad\text{and}\qquad\lambda_{n+1}\geq\lambda_{n+2}\geq\cdots\geq\lambda_{2n}.$
For $\lambda\in\Lambda$, we denote by $\lambda^{\mathrm{antidom}}$ the unique antidominant weight in the dot-orbit $W\cdot\lambda$.
Definition 2.7.
We define the atypicality of a simple module $L=L^{\mathfrak{b}}(\lambda)$ as
[…]
This definition is independent of the choice of $\mathfrak{b}$. If $\mathrm{aty}L=0$, then $L$ is called typical; otherwise, it is called atypical.
Lemma 2.8.
For $\mathfrak{b},\mathfrak{b}^{\prime}\in L(n,n)$ and $\lambda,\lambda^{\prime}\in\Lambda$, the following statements hold:
  1. 1.
    $\operatorname{ch}M^{\mathfrak{b}}(\lambda-\rho^{\mathfrak{b}})=\operatorname{ch}M^{\mathfrak{b}^{\prime}}(\lambda-\rho^{\mathfrak{b}^{\prime}})$;
  2. 2.
    $\dim\operatorname{Hom}(M^{\mathfrak{b}}(\lambda-\rho^{\mathfrak{b}}),M^{\mathfrak{b}^{\prime}}(\lambda-\rho^{\mathfrak{b}^{\prime}}))=1$;
Definition 2.9.
Let $\mathfrak{b}\in L(n,n)$. The category $\mathcal{O}^{\mathfrak{b}}$ is defined as the Serre subcategory of $\mathcal{W}$ generated by $\{L^{\mathfrak{b}}(\lambda)\mid\lambda\in\Lambda\}.$ According to Section 2, as an Serre subcategory, it depends only on $\mathfrak{b}_{\overline{0}}$ (however,the highest weight structure depends on $\mathfrak{b}$).
It is well known that the BGG category $\mathcal{O}$ contains all $\mathfrak{b}$–Verma modules for any $\mathfrak{b}\in L(n,n)$. Similarly, by replacing $\mathfrak{g}$ with $\mathfrak{g}_{\overline{0}}$, we define $\mathcal{O}_{\overline{0}}$ as a full subcategory of $\mathcal{W}_{\overline{0}}$ .
We define the category of modules with an even Verma flag, $\mathcal{F}\!\Delta_{\bar{0}}\subset\mathcal{O}_{\bar{0}}$.
Proposition 2.10 ( [22] ).
For a weight $\lambda$, the following are equivalent:
  1. 1.
    $\lambda$ is antidominant.
  2. 2.
    $L^{()}(\lambda)$ injectively maps to some $M$ s.t. $\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}M\in\mathcal{F}\!\Delta_{\bar{0}}.$
Proof. Note that there exist an injective hom
$L_{\bar{0}}(\lambda)\hookrightarrow\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}L^{()}(\lambda).$ (1)
$\square$
Definition 2.11.
Define a map $(\cdot)^{c}:\mathfrak{gl}(n|n)\to\mathfrak{gl}(n|n)$ on matrix units by
$(e_{ij})^{c}:=\begin{cases}e_{w_{0}(j),\,w_{0}(i)},&\text{if }e_{ij}\in\mathfrak{b}^{(n^{n})}_{\bar{1}},\\[2.0pt] -\,e_{w_{0}(j),\,w_{0}(i)},&\text{otherwise},\end{cases}$
where $w_{0}\in S_{2n}$ is the longest element. Then $(\cdot)^{c}$ extends to a Lie superalgebra automorphism of $\mathfrak{gl}(n|n)$. Moreover,
$(\mathfrak{b}^{\mathrm{st}}_{\bar{0}})^{c}=\mathfrak{b}^{\mathrm{st}}_{\bar{0}},$
hence $(\cdot)^{c}$ induces an involution on $L(n,n)$. This involution corresponds to taking the complement inside the $n\times n$ box. In particular, for $\lambda\in\Lambda$, $M^{()}(\lambda)^{c}$ is an $(n^{n})$-Verma module.
Definition 2.12 ([4, 23]).
Define a map $(\cdot)^{at}:\mathfrak{gl}(n|n)\to\mathfrak{gl}(n|n)$ on matrix units by
$(e_{ij})^{at}:=\begin{cases}e_{2n-j+1,\,2n-i+1},&\text{if }e_{ij}\in\mathfrak{b}^{(n^{n})}_{\bar{1}},\\[2.0pt] -\,e_{2n-j+1,\,2n-i+1},&\text{otherwise}.\end{cases}$
Then $(\cdot)^{at}$ extends to a Lie superalgebra automorphism of $\mathfrak{gl}(n|n)$. Moreover,
$(\mathfrak{b}^{\mathrm{st}}_{\bar{0}})^{at}=\mathfrak{b}^{\mathrm{st}}_{\bar{0}},$
so $(\cdot)^{at}$ induces an involution on $L(n,n)$. This involution corresponds to taking antidiagonal transpose. In particular, we have
$M^{()}(\lambda_{1},\dots,\lambda_{n}\mid\lambda_{n+1},\dots,\lambda_{2n})^{at}\ \cong\ M^{()}\!\bigl(\lambda_{2n},\dots,\lambda_{n+1}\mid\lambda_{n},\dots,\lambda_{1}\bigr).$
3 Brundan-Goodwin functors
Definition 3.1.
Fix the standard basis indices with parity
$|1|=\cdots=|n|=\bar{0},\qquad|n+1|=\cdots=|2n|=\bar{1}.$
Let $\mathfrak{g}=\mathfrak{gl}(n|n)$ with matrix units $e_{i,j}$. Define a $\mathbb{Z}$-grading on $\mathfrak{g}=\mathfrak{gl}(n|n)$ by
$\deg(e_{i,j}):=\begin{cases}j-i,&(1\leq i,j\leq n)\ \text{or}\ (n+1\leq i,j\leq 2n),\\ j-i-n,&(1\leq i\leq n<j\leq 2n),\\ j-i+n,&(n+1\leq i\leq 2n,\ 1\leq j\leq n),\end{cases}$
and set
[…]
We call this the principal good grading.
Definition 3.2.
Let $\mathfrak{g}=\bigoplus_{k\in\mathbb{Z}}\mathfrak{g}(k)$ be the principal good grading above. Set
$\mathfrak{p}_{\mathrm{pg}}:=\bigoplus_{k\geq 0}\mathfrak{g}(k),\qquad\mathfrak{l}_{\mathrm{pg}}:=\mathfrak{g}(0),\qquad\mathfrak{m}_{\mathrm{pg}}:=\bigoplus_{k<0}\mathfrak{g}(k).$
Then $\mathfrak{p}_{\mathrm{pg}}$ is a parabolic subsuperalgebra of $\mathfrak{g}$ with Levi factor $\mathfrak{l}_{\mathrm{pg}}$. Moreover,
$\mathfrak{l}_{\mathrm{pg}}\cong\mathfrak{gl}(1|1)^{\oplus n}.$
Definition 3.3.
The Brundan–Goodwin functor is the exact functor
$\operatorname{BG}:=\operatorname{Ind}_{\mathfrak{p}_{\mathrm{pg}}}^{\mathfrak{g}}\circ\operatorname{Infl}_{\mathfrak{l}_{\mathrm{pg}}}^{\mathfrak{p}_{\mathrm{pg}}}:\ \mathfrak{gl}(1|1)^{\oplus n}\text{-sMod}\longrightarrow\mathfrak{gl}(n|n)\text{-sMod}.$
Recall that $L(1,1)=\{(),(1)\}$ consists of two Borels of $\mathfrak{gl}(1|1)$, namely the distinguished Borel $()=(0)$ and the anti-distinguished Borel $(1)$.
Definition 3.4.
We identify $\gamma=(\gamma_{1},\dots,\gamma_{n})\in\{0,1\}^{n}$ with
$\gamma:=({\bar{\gamma}_{1}})\oplus\cdots\oplus({\bar{\gamma}_{n}})\ \subset\ \mathfrak{gl}(1|1)^{\oplus n}.$
We also write
$\mathbf{o}:=(0,\dots,0)\in\{0,1\}^{n},\qquad\mathbf{i}:=(1,\dots,1)\in\{0,1\}^{n}.$
Definition 3.5 (hypercube Borel subalgebras).
For $\gamma=(\gamma_{1},\dots,\gamma_{n})\in\{0,1\}^{n}$ define a Borel subalgebra $\mathfrak{b}_{\gamma}\subset\mathfrak{gl}(n|n)$ by the partition
$\mathfrak{b}_{\gamma}\ \leftrightarrow\ \bigl(n\!-\!1+\gamma_{n},\ n\!-\!2+\gamma_{n-1},\ \dots,\ 1+\gamma_{2},\ \gamma_{1}\bigr)\ \in L(n,n).$
(Here we use the convention from §2 identifying Borels with partitions in the $n\times n$ rectangle.)
Remark 3.6.
The Borels $\mathfrak{b}_{\mathbf{o}}$ and $\mathfrak{b}_{\mathbf{i}}$ are distinguished in the following sense. Let $\mathfrak{b}=(\beta_{1},\dots,\beta_{n})\in L(n,n)$ with $n\geq\beta_{1}\geq\cdots\geq\beta_{n}\geq 0$, and for $1\leq j\leq n$ set the transpose partition
$\beta^{\prime}_{j}:=\#\{\,r\in\{1,\dots,n\}\mid\beta_{r}\geq j\,\}.$
Then
$\rho_{\bar{0}}=\sum_{i=1}^{n}\frac{n+1-2i}{2}\,\varepsilon_{i}+\sum_{j=1}^{n}\frac{n+1-2j}{2}\,\delta_{j},$
$\rho_{\bar{1}}^{\mathfrak{b}}=\sum_{i=1}^{n}\frac{n-2\beta_{n+1-i}}{2}\,\varepsilon_{i}+\sum_{j=1}^{n}\frac{2\beta^{\prime}_{j}-n}{2}\,\delta_{j},$
and hence
$\rho_{\bar{0}}-\rho_{\bar{1}}^{\mathfrak{b}}=\sum_{i=1}^{n}\frac{2\beta_{n+1-i}-2i+1}{2}\,\varepsilon_{i}+\sum_{j=1}^{n}\frac{2n+1-2j-2\beta^{\prime}_{j}}{2}\,\delta_{j}.$
Moreover,
$\rho^{\mathfrak{b}}=\rho_{\bar{0}}-\rho_{\bar{1}}^{\mathfrak{b}}+\tfrac{1}{2}\,\operatorname{ber}=\sum_{i=1}^{n}\bigl(\beta_{n+1-i}-i+1\bigr)\,\varepsilon_{i}+\sum_{j=1}^{n}\bigl(n-j-\beta^{\prime}_{j}\bigr)\,\delta_{j},$
so
[…]
In the first case $\rho^{\mathfrak{b}}=0$, and in the second case $\rho^{\mathfrak{b}}=\operatorname{ber}$.
Note that if $\lambda=\sum_{i=1}^{m}\lambda_{i}\varepsilon_{i}-\sum_{i=1}^{n}\lambda_{m+i}\delta_{i}$, then
$M^{\mathfrak{b}_{\mathbf{o}}}(\lambda_{1},\dots,\lambda_{m}\mid\lambda_{m+1},\dots,\lambda_{m+n})=M^{\mathfrak{b}_{\mathbf{o}}}(\lambda).$
Moreover, for $\mathfrak{b}\in L(n,n)$, all $\mathfrak{b}$-simple roots are odd if and only if $\mathfrak{b}=\mathfrak{b}_{\mathbf{o}}$ or $\mathfrak{b}=\mathfrak{b}_{\mathbf{i}}$.
Note that $\mathfrak{p}_{\mathrm{pg}}=\mathfrak{b}_{\mathbf{o}}+\mathfrak{b}_{\mathbf{i}}$ and $\mathfrak{m}_{\mathrm{pg}}=\bigoplus_{\alpha\in\Delta_{\bar{0}}^{+}}\mathfrak{g}_{-\alpha}\ \oplus\ \bigoplus_{\alpha\in\Delta_{\bar{1}}^{\mathfrak{b}_{\mathbf{o}}+}\cap\Delta_{\bar{1}}^{\mathfrak{b}_{\mathbf{i}}+}}\mathfrak{g}_{-\alpha}$.
Definition 3.7.
For $\lambda\in\mathfrak{h}^{*}$ we set
$\operatorname{BG}(\lambda):=\operatorname{BG}\bigl(L^{\mathbf{o}}(\lambda)\bigr).$
$\Lambda^{\operatorname{BG}}:=\Bigl\{\ \lambda\in\Lambda\ \Bigm|\ \operatorname{aty}\!\left(L^{()}(\lambda)\right)=\sum_{i=1}^{n}\operatorname{aty}\!\left(L^{()}\!\bigl(\lambda_{i}\,|\,\lambda_{n+i}\bigr)\right)\ \Bigr\}.$
$\Lambda^{\operatorname{maBG}}:=\Bigl\{\ \lambda\in\Lambda\ \Bigm|\sum_{i=1}^{n}\operatorname{aty}\!\left(L^{()}\!\bigl(\lambda_{i}\,|\,\lambda_{n+i}\bigr)\right)=n\ \Bigr\}.$
Remark 3.8 ([4]).
Under the principal Whittaker coinvariants functor $H_{0}$, the module $\operatorname{BG}(\lambda)$ is sent to the corresponding irreducible evaluation module for the (super) Yangian. Moreover,
$H_{0}\operatorname{BG}(\lambda)\ \text{is simple}\ \iff\ \lambda\in\Lambda^{\operatorname{BG}},\qquad\dim H_{0}\operatorname{BG}(\lambda)=1\ \iff\ \lambda\in\Lambda^{\operatorname{maBG}}.$
Proposition 3.9.
Let $\gamma\in\{0,1\}^{n}$ and $\lambda,\mu\in\Lambda$.
  1. 1.
    There is an isomorphism
    $\operatorname{BG}\!\left(M^{\gamma}(\lambda)\right)\ \cong\ M^{\mathfrak{b}_{\gamma}}(\lambda).$
  2. 2.
    There is an isomorphism
    $\operatorname{BG}(\lambda)\ \cong\ \operatorname{Im}\!\left(M^{\mathfrak{b}_{\mathbf{o}}}(\lambda)\xrightarrow{\ \neq 0\ }M^{\mathfrak{b}_{\mathbf{i}}}(\lambda)\right).$
  3. 3.
    We have
    $\operatorname{ch}\operatorname{BG}(\lambda)=\operatorname{ch}L^{\mathbf{o}}(\lambda)\;\prod_{\alpha\in\Delta_{\bar{0}}^{\mathrm{st}+}}\frac{1}{1-e^{-\alpha}}\;\prod_{\gamma\in\Delta_{\bar{1}}^{\mathfrak{b}_{\mathbf{o}}+}\cap\Delta_{\bar{1}}^{\mathfrak{b}_{\mathbf{i}}+}}\left(1+e^{-\gamma}\right).$
    Moreover,
    $\lambda\in\Lambda^{\operatorname{maBG}}$ $\iff\operatorname{ch}L^{\mathbf{o}}(\lambda)=e^{\lambda}$
    $\iff\operatorname{ch}\operatorname{BG}(\lambda)=e^{\lambda}\;\prod_{\alpha\in\Delta_{\bar{0}}^{\mathrm{st}+}}\frac{1}{1-e^{-\alpha}}\;\prod_{\gamma\in\Delta_{\bar{1}}^{\mathfrak{b}_{\mathbf{o}}+}\cap\Delta_{\bar{1}}^{\mathfrak{b}_{\mathbf{i}}+}}\left(1+e^{-\gamma}\right).$
  4. 4.
    Let $B$ be a block of $\mathcal{O}\!\bigl(\mathfrak{gl}(1|1)^{\oplus n}\bigr)$ and let $\Lambda^{B}\subset\Lambda$ be the corresponding set of highest weights. Then $\operatorname{BG}$ restricts to an equivalence
    $B\ \xrightarrow{\ \sim\ }\ \operatorname{Filt}_{\mathcal{O}(\mathfrak{gl}(n|n))}\Bigl\{\operatorname{BG}(\lambda)\ \big|\ \lambda\in\Lambda^{B}\Bigr\}.$
  5. 5.
    $\bigl[M^{\mathbf{o}}(\lambda):L^{\mathbf{o}}(\mu)\bigr]\ =\ \bigl[M^{\mathfrak{b}_{\mathbf{o}}}(\lambda):\operatorname{BG}(\mu)\bigr].$
Proof.
  1. 1.
    This follows from the character identity together with the universal property of induction.
  2. 2.
    This follows from the fact that for $\mathfrak{gl}(1|1)$ the simple module can be realized as the image of a nonzero homomorphism between Verma modules attached to different Borel subalgebras, together with the exactness of $\operatorname{BG}$.
  3. 3.
    This is immediate from the definitions.
  4. 4.
    One checks that for $\lambda,\mu\in\Lambda^{B}$ every morphism
    $\operatorname{BG}(\lambda)\longrightarrow\operatorname{BG}(\mu)$
    is either zero or an isomorphism. Hence $\{\operatorname{BG}(\lambda)\mid\lambda\in\Lambda^{B}\}$ forms a semibrick. By a result of Ringel, the category
    $\operatorname{Filt}_{\mathcal{O}(\mathfrak{gl}(n|n))}\Bigl\{\operatorname{BG}(\lambda)\ \big|\ \lambda\in\Lambda^{B}\Bigr\}$
    is an $\operatorname{Ext}^{1}$-closed abelian subcategory. For details in a more general setting, see [19].
  5. 5.
    This follows from the exactness of $\operatorname{BG}$.
$\square$
Definition 3.10.
Let $\pi:\mathcal{O}\to\bar{\mathcal{O}}$ be the Serre quotient functor by the Serre subcategory generated by non-antidominant simples .
One of the main results of [4] is the following.
Theorem 3.11 ([3, 4]).
  1. 1.
    $\pi\,\operatorname{BG}(\lambda)$ is simple if and only if $\lambda\in\Lambda^{\operatorname{BG}}$.
  2. 2.
    If $\lambda\in\Lambda^{\operatorname{BG}}$, then
    $\pi\,\operatorname{BG}(\lambda)\ \cong\ \pi\,L^{()}\!\bigl((\lambda-\rho)^{\mathrm{antidom}}\bigr).$
  3. 3.
    For $\mathfrak{b}\in L(n,n)$, $w\in W$, and $\lambda\in\Lambda$, the class
    $\bigl[\ \pi\,M^{\mathfrak{b}}(w\lambda-\rho^{\mathfrak{b}})\ \bigr]\ \in\ K_{0}(\bar{\mathcal{O}})$
    is independent of the choice of $w$ and $\mathfrak{b}$.
Theorem 3.12.
If $\lambda\in\Lambda^{\operatorname{BG}}$, then $\operatorname{soc}\operatorname{BG}(\lambda)\cong L^{()}\!\bigl((\lambda-\rho)^{\mathrm{antidom}}\bigr)$.
Proof. Since $\operatorname{BG}(\lambda)\in\mathcal{F}\!\Delta_{\bar{0}}$, Theorem 3.11 implies that $\operatorname{BG}(\lambda)$ has a unique antidominant composition factor. On the other hand, Proposition 2 shows that every simple subquotient of $\operatorname{soc}\operatorname{BG}(\lambda)$ is antidominant. Hence $\operatorname{soc}\operatorname{BG}(\lambda)$ must be simple. $\square$
Theorem 3.13.
Let $\mathfrak{b}\in L(n,n)$ and $\lambda\in\Lambda$. Fix a choice $\lambda^{\operatorname{BG}}\in W\lambda\cap\Lambda^{\operatorname{BG}}$. Then
$\bigoplus_{\nu\in\Lambda^{\mathrm{antidom}}}\bigl(L^{()}(\nu)\bigr)^{\oplus\,\bigl[M^{\mathfrak{b}}(\lambda-\rho^{\mathfrak{b}}):L^{()}(\nu)\bigr]}\ \cong\ \bigoplus_{\mu\in\Lambda}\bigl(L^{()}\!\bigl((\mu-\rho)^{\mathrm{antidom}}\bigr)\bigr)^{\oplus\,\bigl[M^{\mathbf{o}}(\lambda^{\operatorname{BG}}):L^{\mathbf{o}}(\mu)\bigr]}.$
Proof. In $\bar{\mathcal{O}}$ we have
$\bigl[\pi\,M^{\mathfrak{b}}(\lambda-\rho^{\mathfrak{b}})\bigr]$ $=\bigl[\pi\,M^{\mathfrak{b}_{\mathbf{o}}}(\lambda^{\operatorname{BG}})\bigr]$
$=\sum_{\mu\in\Lambda}\bigl[\pi\,\operatorname{BG}(\mu)\bigr]^{\oplus\,\bigl[M^{\mathfrak{b}_{\mathbf{o}}}(\lambda^{\operatorname{BG}}):\operatorname{BG}(\mu)\bigr]}$
$=\sum_{\mu\in\Lambda}\bigl[\pi\,L^{()}\!\bigl((\mu-\rho)^{\mathrm{antidom}}\bigr)\bigr]^{\oplus\,\bigl[M^{\mathbf{o}}(\lambda^{\operatorname{BG}}):L^{\mathbf{o}}(\mu)\bigr]}.$
Note that the multiplicity $\bigl[M^{\mathbf{o}}(\lambda^{\operatorname{BG}}):L^{\mathbf{o}}(\mu)\bigr]$ does not depend on the choice of $\lambda^{\operatorname{BG}}$. Note also that if $\bigl[M^{\mathbf{o}}(\lambda^{\operatorname{BG}}):L^{\mathbf{o}}(\mu)\bigr]\neq 0$, then $\mu\in\Lambda^{\operatorname{BG}}$. $\square$
The socle of a $\mathfrak{b}$-Verma supermodule is known when $\mathfrak{b}=()$ or $\mathfrak{b}=(n^{n})$ by [7]; however, it remains an open problem in general. The following consequence of the previous theorem will be fundamental for this problem.
Corollary 3.14.
Let $\mathfrak{b}\in L(n,n)$ and $\lambda\in\Lambda$. Fix a choice $\lambda^{\operatorname{BG}}\in W\lambda\cap\Lambda^{\operatorname{BG}}$. Then $\operatorname{soc}M^{\mathfrak{b}}(\lambda-\rho^{\mathfrak{b}})$ is a direct summand of
$\bigoplus_{\mu\in\Lambda}L^{()}\!\bigl((\mu-\rho)^{\mathrm{antidom}}\bigr)^{\oplus\,\bigl[M^{\mathbf{o}}(\lambda^{\operatorname{BG}}):L^{\mathbf{o}}(\mu)\bigr]}.$
4 Duflo-Serganova functors
Definition 4.1.
Define a subset $X\subseteq\mathfrak{g}_{\overline{1}}$ by:
$X:=\{x\in\mathfrak{g}_{\overline{1}}\mid[x,x]=0\}.$
Forx in X , $M\in\mathfrak{g}\text{-sMod}$, we define a supervector space $DS_{x}M$ by:
$DS_{x}M:=\ker x_{M}/\operatorname{Im}x_{M},$
where $x_{M}$ denotes the action of $x$ on $M$.
Proposition 4.2 ([14]).
Let $x\in X$. For the adjoint module $\mathfrak{g}\in\mathfrak{g}\text{-sMod}$, the module
$\mathfrak{g}_{x}:=\operatorname{DS}_{x}(\mathfrak{g})$
naturally inherits the structure of a basic Lie superalgebra. Consequently, there is a natural symmetric monoidal $k$-linear functor
$\operatorname{DS}_{x}:\ \mathfrak{g}\text{-sMod}\longrightarrow\mathfrak{g}_{x}\text{-sMod},$
called the Duflo–Serganova functor.
Lemma 4.3 (Hinich’s Lemma [14]).
Let $x\in X$. Given a short exact sequence
$0\rightarrow L\rightarrow M\rightarrow N\rightarrow 0$
in $\mathfrak{g}\text{-sMod}$, there exists $E\in\mathfrak{g}_{x}\text{-sMod}$ such that the following sequence is exact in $\mathfrak{g}_{x}\text{-sMod}$:
$0\to\Pi E\to DS_{x}L\to DS_{x}M\to DS_{x}N\to E\to 0.$
Lemma 4.4 ([14], Lemma 2.20).
Let $x\in X$ and $M\in\mathfrak{g}_{\bar{0}}\text{-sMod}$. Then
$\operatorname{DS}_{x}\!\left(\operatorname{Ind}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}M\right)=0.$
In particular, if $N\in\mathcal{O}$ belongs to a typical block, then $\operatorname{DS}_{x}(N)=0$.
Proposition 4.5.
Let $\mathfrak{g}=\mathfrak{gl}(n|n)$ and let $x=e_{i,j}\in\mathfrak{g}_{\bar{1}}$ be an odd root vector. Then $\mathfrak{g}_{x}\cong\mathfrak{gl}(n-1|n-1)$ can be realized as a Lie subsuperalgebra of $\mathfrak{g}$. Moreover, for every Borel subalgebra $\mathfrak{b}\subset\mathfrak{g}$ we have
$\mathfrak{b}_{x}\cong\mathfrak{b}\cap\mathfrak{g}_{x},$
and $\mathfrak{b}_{x}$ is a Borel subalgebra of $\mathfrak{g}_{x}$.
Example 4.6.
Let $\mathfrak{g}=\mathfrak{gl}(1|1)$ and let $x=e_{1,2}\in\mathfrak{g}_{\bar{1}}$. For an atypical weight $\lambda$ one has
$\operatorname{DS}_{x}\!\left(L^{()}(\lambda)\right)\cong\Pi^{\operatorname{par}(\lambda)}\,k,\qquad\operatorname{DS}_{x}\!\left(M^{()}(\lambda)\right)\cong k\oplus\Pi k,$
and moreover
$\operatorname{DS}_{x}\!\left(M^{(1)}(\lambda)\right)\cong 0,\qquad\operatorname{DS}_{x}\!\left(P^{()}(\lambda)\right)\cong 0.$
Here $\operatorname{DS}_{x}$ takes values in $\mathfrak{g}_{x}\text{-sMod}$, and in this case $\mathfrak{g}_{x}\cong\mathfrak{gl}(0|0)$, hence $\mathfrak{g}_{x}\text{-sMod}\cong\operatorname{sVec}$.
Let $\alpha:=\varepsilon_{1}-\delta_{1}$. There is a short exact sequence
$0\to L^{()}(\lambda-\alpha)\to M^{()}(\lambda)\to L^{()}(\lambda)\to 0.$
Applying $\operatorname{DS}_{x}$ gives an exact sequence
$0\to 0\to\Pi\,\Pi^{\operatorname{par}(\lambda)}k\to k\oplus\Pi k\to\Pi^{\operatorname{par}(\lambda)}k\to 0\to 0.$
Similarly, there is a short exact sequence
$0\to L^{()}(\lambda+\alpha)\to M^{(1)}(\lambda)\to L^{()}(\lambda)\to 0,$
and applying $\operatorname{DS}_{x}$ gives an exact sequence
$0\to\Pi\,\Pi^{\operatorname{par}(\lambda)}k\to\Pi\,\Pi^{\operatorname{par}(\lambda)}k\to 0\to\Pi^{\operatorname{par}(\lambda)}k\to\Pi^{\operatorname{par}(\lambda)}k\to 0.$
Example 4.7.
Let $\mathfrak{g}=\mathfrak{gl}(1|1)\oplus\mathfrak{gl}(n|n)$ and let $x=e_{1,2}\in\mathfrak{gl}(1|1)\subset\mathfrak{g}$. Then for $M\in\mathfrak{gl}(1|1)\text{-sMod}$ and $N\in\mathfrak{gl}(n|n)\text{-sMod}$ there is a natural isomorphism
$\operatorname{DS}_{x}(M\boxtimes N)\ \cong\ \operatorname{DS}_{x}(M)\boxtimes N.$
Here $\operatorname{DS}_{x}$ takes values in $\mathfrak{g}_{x}\text{-sMod}$ with $\mathfrak{g}_{x}\cong\mathfrak{gl}(n|n)$.
5 Main results
Theorem 5.1.
Let $\mathfrak{g}=\mathfrak{gl}(n|n)$ with the standard parity $|1|=\cdots=|n|=\bar{0}$ and $|n+1|=\cdots=|2n|=\bar{1}$. Set $I=\{1,n+1\}$ and $J=\{2,\dots,n\}\cup\{n+2,\dots,2n\}$, and define
$\mathfrak{l}_{IJ}:=\mathfrak{gl}(I)\oplus\mathfrak{gl}(J)\cong\mathfrak{gl}(1|1)\oplus\mathfrak{gl}(n-1|n-1),\qquad\mathfrak{p}_{IJ}:=\mathfrak{l}_{IJ}\oplus\mathfrak{u}_{IJ},$
[…]
$\operatorname{DS}^{\mathfrak{g}}_{e_{1,n+1}}\!\left(\operatorname{Ind}_{\mathfrak{p}_{IJ}}^{\mathfrak{g}}\,\operatorname{Infl}_{\mathfrak{l}_{IJ}}^{\mathfrak{p}_{IJ}}(M)\right)\cong\operatorname{DS}^{\mathfrak{l}_{IJ}}_{e_{1,n+1}}(M).$
Equivalently, as functors $\mathfrak{l}_{IJ}\text{-sMod}\to\mathfrak{g}_{e_{1,n+1}}\text{-sMod}$,
$\operatorname{DS}^{\mathfrak{g}}_{e_{1,n+1}}\circ\operatorname{Ind}_{\mathfrak{p}_{IJ}}^{\mathfrak{g}}\circ\operatorname{Infl}_{\mathfrak{l}_{IJ}}^{\mathfrak{p}_{IJ}}\ \cong\ \operatorname{DS}^{\mathfrak{l}_{IJ}}_{e_{1,n+1}}.$
Proof. Let
$V:=\operatorname{Ind}_{\mathfrak{p}_{IJ}}^{\mathfrak{g}}\operatorname{Infl}_{\mathfrak{l}_{IJ}}^{\mathfrak{p}_{IJ}}(M).$
Define the differential $d$ on $V$ by
$d(v):=e_{1,n+1}\cdot v\qquad(v\in V).$
Since $e_{1,n+1}$ is odd and $e_{1,n+1},e_{1,n+1}]=0$, we have $d^{2}=0$ and
$H(V,d)=\ker(d)/\mathrm{im}(d)=\operatorname{DS}^{\mathfrak{g}}_{e_{1,n+1}}(V).$
Step 1: identify $V$ with $S(\mathfrak{u}^{-})\otimes M$ and write $d$. By PBW and the vector space decomposition $\mathfrak{g}=\mathfrak{u}^{-}\oplus\mathfrak{p}_{IJ}$, multiplication induces a linear isomorphism
$V\ \cong\ U(\mathfrak{u}^{-})\otimes M.$
A direct check from the $\mathfrak{gl}(n|n)$ bracket shows $\mathfrak{u}^{-}$ is abelian, hence
$U(\mathfrak{u}^{-})\ \cong\ S(\mathfrak{u}^{-}).$
We thus view $V$ as $S(\mathfrak{u}^{-})\otimes M$.
For homogeneous $u\in S(\mathfrak{u}^{-})$ and $m\in M$, we have
$d(u\otimes m)=(e_{1,n+1}u)\otimes m.$
Super-commuting $e_{1,n+1}$ to the right and using that the tensor product is over $U(\mathfrak{p}_{IJ})$ (so $(ue_{1,n+1})\otimes m=u\otimes(e_{1,n+1}m)$), we obtain
$d(u\otimes m)=[e_{1,n+1},u]\otimes m\;+\;(-1)^{|u|}\,u\otimes(e_{1,n+1}m).$ (2)
Since $\mathfrak{u}^{-}$ is abelian, the map $u\mapsto[e_{1,n+1},u]$ is an odd super-derivation of $S(\mathfrak{u}^{-})$.
Step 2: a contraction on $(S(\mathfrak{u}^{-}),[e_{1,n+1},\cdot])$. Set
$\delta:=[e_{1,n+1},\,\cdot\,]:S(\mathfrak{u}^{-})\to S(\mathfrak{u}^{-}).$
For each $i\in J$, the standard superbracket relation gives
$\delta(e_{i,1})=[e_{1,n+1},e_{i,1}]=-(-1)^{|i|}\,e_{i,n+1},\qquad\delta(e_{i,n+1})=[e_{1,n+1},e_{i,n+1}]=0.$ (3)
[…]
$h(1)=0,\qquad h(e_{i,1})=0,\qquad h(e_{i,n+1})=-(-1)^{|i|}e_{i,1}\quad(i\in J),$
and extend $h$ to all of $S(\mathfrak{u}^{-})$ as an odd super-derivation:
$h(uv)=h(u)v+(-1)^{|u|}u\,h(v)\qquad(u,v\ \text{homogeneous}).$
Define also an even derivation $D:S(\mathfrak{u}^{-})\to S(\mathfrak{u}^{-})$ by
$D(1)=0,\qquad D(e_{i,1})=e_{i,1},\qquad D(e_{i,n+1})=e_{i,n+1}\quad(i\in J),$
and uniquely extend by the (even) Leibniz rule $D(uv)=D(u)v+uD(v)$. Thus $D$ acts on a homogeneous monomial of total degree $k$ as multiplication by $k$. We claim that
$\delta h+h\delta=D\qquad\text{on }S(\mathfrak{u}^{-}).$
Indeed, first note that $(\delta h+h\delta)(1)=0=D(1)$. Next, for $e_{i,1}$ we have $h(e_{i,1})=0$, and using (3) we compute
$(\delta h+h\delta)(e_{i,1})=h(\delta(e_{i,1}))=h\bigl(-(-1)^{|i|}e_{i,n+1}\bigr)=-(-1)^{|i|}\,h(e_{i,n+1})=e_{i,1}=D(e_{i,1}).$
For $e_{i,n+1}$ we have $\delta(e_{i,n+1})=0$, hence
$(\delta h+h\delta)(e_{i,n+1})=\delta(h(e_{i,n+1}))=\delta\bigl(-(-1)^{|i|}e_{i,1}\bigr)=-(-1)^{|i|}\,\delta(e_{i,1})=e_{i,n+1}=D(e_{i,n+1}).$
Finally, $\delta$ and $h$ are odd super-derivations, so $\delta h+h\delta$ is an even derivation; by construction $D$ is also an even derivation. Since these two even derivations agree on the algebra generators $\{e_{i,1},e_{i,n+1}\mid i\in J\}$, they agree on all of $S(\mathfrak{u}^{-})$ by the Leibniz rule. Hence $\delta h+h\delta=D$ on $S(\mathfrak{u}^{-})$.
Decompose $S(\mathfrak{u}^{-})=\bigoplus_{k\geq 0}S_{k}$ by total degree, where $S_{k}$ is spanned by supermonomials of degree $k$. Then $D|_{S_{k}}=k\cdot\operatorname{id}$, so $D$ is invertible on
[…]
Define an odd map $s:S(\mathfrak{u}^{-})\to S(\mathfrak{u}^{-})$ by setting $s|_{S(\mathfrak{u}^{-})_{>0}}:=D^{-1}\circ h$ and $s|_{S_{0}}:=0$ (equivalently, $s(1)=0$). Then, using $\delta h+h\delta=D$, we have on $S(\mathfrak{u}^{-})_{>0}$:
$(\delta s+s\delta)=\delta(D^{-1}h)+(D^{-1}h)\delta=D^{-1}(\delta h+h\delta)=D^{-1}D=\operatorname{id}.$
Since $(\delta s+s\delta)(1)=0$, this can be written on all of $S(\mathfrak{u}^{-})$ as
$\delta s+s\delta=\operatorname{id}-\pi,$
[…]
Step 3: tensor with $M$ and conclude. Write (2) as $d=d_{1}+d_{2}$ where
$d_{1}(u\otimes m):=\delta(u)\otimes m,\qquad d_{2}(u\otimes m):=(-1)^{|u|}u\otimes(e_{1,n+1}m).$
From Step 2 we have an odd map $s:S(\mathfrak{u}^{-})\to S(\mathfrak{u}^{-})$ such that
$\delta s+s\delta=\mathrm{id}-\pi.$
Tensoring with $\operatorname{id}_{M}$ gives
$d_{1}\,(s\otimes\operatorname{id}_{M})+(s\otimes\operatorname{id}_{M})\,d_{1}=\operatorname{id}_{S(\mathfrak{u}^{-})\otimes M}-(\pi\otimes\operatorname{id}_{M}).$
Claim. We have the odd–odd anti-commutation relation
$d_{2}\,(s\otimes\operatorname{id}_{M})+(s\otimes\operatorname{id}_{M})\,d_{2}=0.$
Indeed, for homogeneous $u\in S(\mathfrak{u}^{-})$,
$(s\otimes\operatorname{id}_{M})\,d_{2}(u\otimes m)=(-1)^{|u|}s(u)\otimes(e_{1,n+1}m),$
while
$d_{2}\,(s\otimes\operatorname{id}_{M})(u\otimes m)=d_{2}(s(u)\otimes m)=(-1)^{|s(u)|}s(u)\otimes(e_{1,n+1}m)=-(-1)^{|u|}s(u)\otimes(e_{1,n+1}m),$
since $|s(u)|\equiv|u|+1\pmod{2}$ (because $s$ is odd). Hence the two terms sum to 0.
Consequently,
$d\,(s\otimes\operatorname{id}_{M})+(s\otimes\operatorname{id}_{M})\,d=\operatorname{id}_{S(\mathfrak{u}^{-})\otimes M}-(\pi\otimes\operatorname{id}_{M}).$
Hence $\ker(\pi\otimes\operatorname{id}_{M})$ is contractible, and the inclusion
[…]
is a homotopy equivalence with homotopy inverse $\pi\otimes\operatorname{id}_{M}$.
[…]
$H(V,d)\cong\operatorname{DS}^{\mathfrak{l}_{IJ}}_{e_{1,n+1}}(M),$
naturally in $M$. $\square$
Definition 5.2.
Let $\mathfrak{b}_{1|1}=(\beta_{n})\in L(1,1)$ and $\mathfrak{b}_{n-1|n-1}=(\beta_{1},\dots,\beta_{n-1},\beta_{n}=1)\in L(n-1,n-1)$. Define
$\mathfrak{b}_{1|1}\star\mathfrak{b}_{n-1|n-1}:=(\beta_{1}+1,\dots,\beta_{n-1}+1,\beta_{n})\ \in L(n,n).$
Note that $(\mathfrak{b}_{1|1}\star\mathfrak{b}_{n-1|n-1})_{e_{1,n+1}}\cong\mathfrak{b}_{n-1|n-1}.$
Definition 5.3.
For $\lambda\in\Lambda_{n|n}$ we write $\lambda|_{n-1|n-1}\in\Lambda_{n-1|n-1}$ and $\lambda|_{1|1}\in\Lambda_{1|1}$ for the restrictions corresponding to the Levi embedding
$\mathfrak{gl}(1|1)\oplus\mathfrak{gl}(n-1|n-1)\hookrightarrow\mathfrak{gl}(n|n).$
If $\lambda\in\Lambda^{\operatorname{BG}}_{n|n}$ (resp. $\lambda\in\Lambda^{\operatorname{maBG}}_{n|n}$), then
$\lambda|_{n-1|n-1}\in\Lambda^{\operatorname{BG}}_{n-1|n-1}\qquad\text{(resp.\ }\lambda|_{n-1|n-1}\in\Lambda^{\operatorname{maBG}}_{n-1|n-1}\text{)}.$
Lemma 5.4.
Let $\mathfrak{b}_{1|1}\in L(1,1)$, $\mathfrak{b}_{n-1|n-1}\in L(n-1,n-1)$, and $\lambda\in\mathfrak{h}^{*}$ with restrictions $\lambda|_{1|1}$ and $\lambda|_{n-1|n-1}$ as in Remark 5. Then there is an isomorphism
$\mathrm{Ind}_{\mathfrak{p}_{IJ}}^{\mathfrak{g}}\mathrm{Infl}_{\mathfrak{l}_{IJ}}^{\mathfrak{p}_{IJ}}\Bigl(M^{\mathfrak{b}_{1|1}}\bigl(\lambda|_{1|1}\bigr)\ \boxtimes\ M^{\mathfrak{b}_{n-1|n-1}}\bigl(\lambda|_{n-1|n-1}\bigr)\Bigr)\ \cong\ M^{\mathfrak{b}_{1|1}\star\mathfrak{b}_{n-1|n-1}}(\lambda).$
Proof. The two modules have the same character and are both generated by a highest weight vector of the same highest weight. Hence they are isomorphic by the universlity of right hand side. $\square$
Example 5.5.
In $\mathfrak{gl}(3|3)$, every Borel subalgebra in $L(3,3)$ except
[…]
can be written in the form $\mathfrak{b}_{1|1}\star\mathfrak{b}_{2|2}$. By applying the Lie superalgebra automorphism $(\cdot)^{at}$ and arguing similarly, we also obtain that the $(2)$-, $(3)$-, $(2^{3})$-, and $(32^{2})$-Verma modules can be realized via such a parabolic induction.
Lemma 5.6.
Let $M\in\mathcal{O}_{\mathfrak{gl}(1|1)}$ and $N\in\mathcal{O}_{\mathfrak{gl}(1|1)^{\oplus(n-1)}}$. Then there is a natural isomorphism
$\operatorname{Ind}_{\mathfrak{p}_{IJ}}^{\mathfrak{gl}(n|n)}\operatorname{Infl}_{\mathfrak{l}_{IJ}}^{\mathfrak{p}_{IJ}}\Bigl(M\boxtimes\operatorname{BG}_{n-1|n-1}(N)\Bigr)\ \cong\ \operatorname{BG}_{n|n}(M\boxtimes N).$
Proof. It is enough to show that for every $\lambda\in\Lambda_{n|n}$ there is an isomorphism
$\operatorname{Ind}_{\mathfrak{p}_{IJ}}^{\mathfrak{gl}(n|n)}\operatorname{Infl}_{\mathfrak{l}_{IJ}}^{\mathfrak{p}_{IJ}}\Bigl(L^{()}(\lambda|_{1|1})\boxtimes\operatorname{BG}_{n-1|n-1}(\lambda|_{n-1|n-1})\Bigr)\ \cong\ \operatorname{BG}_{n|n}(\lambda).$
Both sides have the same character and are generated by a highest weight vector of highest weight $\lambda$.
To make this explicit, define a subsuperalgebra
$\mathfrak{b}_{\lambda}:=\mathfrak{b}_{\mathbf{o}}\ \oplus\ \bigoplus_{\begin{subarray}{c}\alpha\in\Delta_{\bar{1}}^{\mathfrak{b}_{\mathbf{i}}+}\setminus\Delta_{\bar{1}}^{\mathfrak{b}_{\mathbf{o}}+}\\ (\lambda,\alpha)=0\end{subarray}}\mathfrak{g}_{\alpha}\ \subset\ \mathfrak{g},$
then $\mathfrak{b}_{\lambda}$have an one-dimensional module $k_{\lambda}$ so that $\mathfrak{h}$ acts by $\lambda$ . Then $\operatorname{BG}(\lambda)$ is isomorphic to the induced module
$\operatorname{BG}(\lambda)\ \cong\ \operatorname{Ind}_{\mathfrak{b}_{\lambda}}^{\mathfrak{g}}k_{\lambda}.$
In particular, $\operatorname{BG}(\lambda)$ is characterized by the universal property of induction from $\mathfrak{b}_{\lambda}$. Applying this universal property, the highest weight vector on the left-hand side yields a nonzero homomorphism from the left-hand side to $\operatorname{BG}(\lambda)$, hence a surjection. $\square$
Theorem 5.7.
For $\lambda\in\Lambda$ one has
$\operatorname{DS}_{e_{1,n+1}}\Bigl(M^{()\star\mathfrak{b}_{n-1|n-1}}(\lambda)\Bigr)\ \cong\ \begin{cases}M^{\mathfrak{b}_{n-1|n-1}}\bigl(\lambda|_{n-1|n-1}\bigr)\oplus\Pi\,M^{\mathfrak{b}_{n-1|n-1}}\bigl(\lambda|_{n-1|n-1}\bigr),&\lambda|_{1|1}\ \text{atypical},\\ 0,&\lambda|_{1|1}\ \text{typical}.\end{cases}$
In particular,
$\operatorname{DS}_{e_{1,n+1}}\Bigl(M^{\mathfrak{b}_{\mathbf{o}}}(\lambda)\Bigr)\ \cong\ \begin{cases}M^{\mathfrak{b}_{\mathbf{o}}}\bigl(\lambda|_{n-1|n-1}\bigr)\oplus\Pi\,M^{\mathfrak{b}_{\mathbf{o}}}\bigl(\lambda|_{n-1|n-1}\bigr),&\lambda|_{1|1}\ \text{atypical},\\ 0,&\lambda|_{1|1}\ \text{typical}.\end{cases}$
Proof. We compute
$\operatorname{DS}_{e_{1,n+1}}\Bigl(M^{()\star\mathfrak{b}_{n-1|n-1}}(\lambda)\Bigr)$ $\cong\operatorname{DS}_{e_{1,n+1}}\operatorname{Ind}_{\mathfrak{p}_{IJ}}^{\mathfrak{g}}\operatorname{Infl}_{\mathfrak{l}_{IJ}}^{\mathfrak{p}_{IJ}}\Bigl(M^{\mathfrak{b}_{1|1}}\bigl(\lambda|_{1|1}\bigr)\boxtimes M^{\mathfrak{b}_{n-1|n-1}}\bigl(\lambda|_{n-1|n-1}\bigr)\Bigr)$
$\cong\operatorname{DS}^{\mathfrak{l}_{IJ}}_{e_{1,n+1}}\Bigl(M^{\mathfrak{b}_{1|1}}\bigl(\lambda|_{1|1}\bigr)\boxtimes M^{\mathfrak{b}_{n-1|n-1}}\bigl(\lambda|_{n-1|n-1}\bigr)\Bigr)$
$\cong\Bigl(\operatorname{DS}_{e_{1,n+1}}M^{\mathfrak{b}_{1|1}}\bigl(\lambda|_{1|1}\bigr)\Bigr)\boxtimes M^{\mathfrak{b}_{n-1|n-1}}\bigl(\lambda|_{n-1|n-1}\bigr).$
Here the first isomorphism follows from Lemma 5, the second from Theorem 5.1, and the third from Example 4. Now apply Example 4 to $\operatorname{DS}_{e_{1,n+1}}M^{\mathfrak{b}_{1|1}}\bigl(\lambda|_{1|1}\bigr)$ to obtain the stated cases. $\square$
Theorem 5.8.
Let $M\in\mathcal{O}_{\mathfrak{gl}(1|1)}$ and $N\in\mathcal{O}_{\mathfrak{gl}(1|1)^{\oplus(n-1)}}$. Then there is a natural isomorphism
$\operatorname{DS}_{e_{1,n+1}}\bigl(\operatorname{BG}_{n|n}(M\boxtimes N)\bigr)\ \cong\ \bigl(\operatorname{DS}_{e_{1,2}}M\bigr)\boxtimes\operatorname{BG}_{n-1|n-1}(N),$
where on the right-hand side we regard $\operatorname{DS}_{e_{1,2}}M$ as an object of $\mathfrak{gl}(0|0)\text{-sMod}\cong\operatorname{sVec}$.
Proof. We compute
$\operatorname{DS}_{e_{1,n+1}}\bigl(\operatorname{BG}_{n|n}(M\boxtimes N)\bigr)$ $\cong\operatorname{DS}_{e_{1,n+1}}\operatorname{Ind}_{\mathfrak{p}_{IJ}}^{\mathfrak{gl}(n|n)}\operatorname{Infl}_{\mathfrak{l}_{IJ}}^{\mathfrak{p}_{IJ}}\Bigl(M\boxtimes\operatorname{BG}_{n-1|n-1}(N)\Bigr)$
$\cong\operatorname{DS}^{\mathfrak{l}_{IJ}}_{e_{1,n+1}}\Bigl(M\boxtimes\operatorname{BG}_{n-1|n-1}(N)\Bigr)$
$\cong\Bigl(\operatorname{DS}_{e_{1,2}}M\Bigr)\boxtimes\operatorname{BG}_{n-1|n-1}(N).$
Here the first isomorphism follows from Lemma 5, the second from Theorem 5.1, and the third from Example 4. $\square$
Theorem 5.9.
If $\lambda\in\Lambda^{\operatorname{maBG}}$, then
$\operatorname{DS}_{e_{1,n+1}}\bigl(\operatorname{BG}_{n|n}(\lambda)\bigr)\ \cong\ \Pi^{\operatorname{par}(\lambda)}\,\operatorname{BG}_{n-1|n-1}\bigl(\lambda|_{n-1|n-1}\bigr).$
Proof. It follows from Example 4 and Theorem 5.8. $\square$
Remark 5.10.
By applying the Lie superalgebra automorphisms $(\cdot)^{c}$ and $(\cdot)^{\operatorname{at}}$, one obtains analogous versions of the above results.
6 $\mathfrak{gl}(2|2)$-examles
Example 6.1.
In the maximal atypical block of $\mathfrak{g}=\mathfrak{gl}(2|2)$, let us check that our results are compatible with Hinich’s lemma. One has
$\lambda\in\Lambda^{\operatorname{maBG}}\iff\lambda=(a,b\mid a,b)\quad\text{for some }a,b\in\mathbb{Z}.$
Note that $\mathfrak{b}_{\mathbf{o}}=(1)$ for $\mathfrak{gl}(2|2)$, while $\mathfrak{b}_{\mathbf{o}}=()$ for $\mathfrak{gl}(1|1)$.
(1) There is a short exact sequence
$0\to\operatorname{BG}(a-1,b\mid a-1,b)\to\operatorname{BG}\bigl(M^{()}(a\mid a)\boxtimes L^{()}(b\mid b)\bigr)\to\operatorname{BG}(a,b\mid a,b)\to 0.$
Applying $\operatorname{DS}_{e_{1,3}}$ gives an exact sequence
$0\to 0\to\Pi\,\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b\mid b)\to L^{()}(b\mid b)\oplus\Pi L^{()}(b\mid b)\to\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b\mid b)\to 0\to 0.$
(2) There is a short exact sequence
$0\to\operatorname{BG}(a,b-1\mid a,b-1)\to\operatorname{BG}\bigl(L^{()}(a\mid a)\boxtimes M^{()}(b\mid b)\bigr)\to\operatorname{BG}(a,b\mid a,b)\to 0.$
Applying $\operatorname{DS}_{e_{1,3}}$ gives an exact sequence
$0\to 0\to\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b-1\mid b-1)\to\Pi^{\operatorname{par}(\lambda|_{1|1})}M^{()}(b\mid b)\to\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b\mid b)\to 0\to 0.$
(3) There is a short exact sequence
$0\to\operatorname{BG}\bigl(M^{()}(a\mid a)\boxtimes L^{()}(b-1\mid b-1)\bigr)\to M^{(1)}(a,b\mid a,b)\to\operatorname{BG}\bigl(M^{()}(a\mid a)\boxtimes L^{()}(b\mid b)\bigr)\to 0.$
Applying $\operatorname{DS}_{e_{1,3}}$ gives an exact sequence
$0\to 0\to\;$ $L^{()}(b-1\mid b-1)\oplus\Pi L^{()}(b-1\mid b-1)\to$
$M^{()}(b\mid b)\oplus\Pi M^{()}(b-1\mid b-1)\to L^{()}(b\mid b)\oplus\Pi L^{()}(b\mid b)\to 0\to 0.$
(4) There is a short exact sequence
$0\to\operatorname{BG}\bigl(L^{()}(a-1\mid a-1)\boxtimes M^{()}(b\mid b)\bigr)\to M^{(1)}(a,b\mid a,b)\to\operatorname{BG}\bigl(L^{()}(a\mid a)\boxtimes M^{()}(b\mid b)\bigr)\to 0.$
Applying $\operatorname{DS}_{e_{1,3}}$ gives an exact sequence
$0\to 0\to\Pi\,\Pi^{\operatorname{par}(\lambda|_{1|1})}M^{()}(b\mid b)\to M^{()}(b\mid b)\oplus\Pi M^{()}(b\mid b)\to\Pi^{\operatorname{par}(\lambda|_{1|1})}M^{()}(b\mid b)\to 0\to 0.$
(5) There is a short exact sequence
$0\to\operatorname{BG}(a+1,b\mid a+1,b)\to\operatorname{BG}\bigl(M^{(1)}(a+1\mid a+1)\boxtimes L^{()}(b\mid b)\bigr)\to\operatorname{BG}(a,b\mid a,b)\to 0.$
Applying $\operatorname{DS}_{e_{1,3}}$ gives an exact sequence
$0\to 0\to\Pi\,\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b\mid b)\to 0\to\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b\mid b)\to 0\to 0.$
(6) There is a short exact sequence
$0\to\operatorname{BG}(a,b+1\mid a,b+1)\to\operatorname{BG}\bigl(L^{()}(a\mid a)\boxtimes M^{(1)}(b+1\mid b+1)\bigr)\to\operatorname{BG}(a,b\mid a,b)\to 0.$
Applying $\operatorname{DS}_{e_{1,3}}$ gives an exact sequence
$0\to 0\to\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b+1\mid b+1)\to\Pi^{\operatorname{par}(\lambda|_{1|1})}M^{(1)}(b+1\mid b+1)\to\Pi^{\operatorname{par}(\lambda|_{1|1})}L^{()}(b\mid b)\to 0\to 0.$
(7) There is a short exact sequence
$0\to\operatorname{BG}\bigl(M^{()}(a\mid a)\boxtimes L^{()}(b+1\mid b+1)\bigr)\to M^{(2)}(a,b+1\mid a,b+1)\to\operatorname{BG}\bigl(M^{()}(a\mid a)\boxtimes L^{()}(b\mid b)\bigr)\to 0.$
Applying $\operatorname{DS}_{e_{1,3}}$ gives an exact sequence
$0\to 0\to\;$ $L^{()}(b+1\mid b+1)\oplus\Pi L^{()}(b+1\mid b+1)\to$
$M^{(1)}(b+1\mid b+1)\oplus\Pi M^{(1)}(b+1\mid b+1)\to L^{()}(b\mid b)\oplus\Pi L^{()}(b\mid b)\to 0\to 0.$
(8) There is a short exact sequence
$0\to\operatorname{BG}\bigl(L^{()}(a+1\mid a+1)\boxtimes M^{()}(b\mid b)\bigr)\to M^{(1^{2})}(a+1,b\mid a+1,b)\to\operatorname{BG}\bigl(L^{()}(a\mid a)\boxtimes M^{()}(b\mid b)\bigr)\to 0.$
Applying $\operatorname{DS}_{e_{1,3}}$ gives an exact sequence
$0\to 0\to\Pi\,\Pi^{\operatorname{par}(\lambda|_{1|1})}M^{()}(b\mid b)\to 0\to\Pi^{\operatorname{par}(\lambda|_{1|1})}M^{()}(b\mid b)\to 0\to 0.$
Example 6.2.
Let $\mathfrak{g}=\mathfrak{gl}(2|2)$ and $\lambda\in\Lambda^{\operatorname{maBG}}$. Then $\operatorname{BG}(\lambda)$ admits a PBW-type basis consisting of vectors of the form
$e_{2,1}^{a_{2,1}}\,e_{2,3}^{a_{2,3}}\,e_{4,1}^{a_{4,1}}\,e_{4,3}^{a_{4,3}}\,v,$
where $v$ is a highest weight vector, $a_{2,1},a_{4,3}\in\mathbb{Z}_{\geq 0}$, and $a_{2,3},a_{4,1}\in\mathbb{Z}/2\mathbb{Z}$.
Then
$e_{1,3}\,e_{2,1}^{a_{2,1}}e_{2,3}^{a_{2,3}}e_{4,1}^{a_{4,1}}e_{4,3}^{a_{4,3}}v=\begin{cases}-\,a_{2,1}\,e_{2,1}^{a_{2,1}-1}\,e_{2,3}\,e_{4,3}^{a_{4,3}}v,&(a_{2,3},a_{4,1})=(0,0),\\[4.0pt] 0,&(a_{2,3},a_{4,1})=(1,0),\\[4.0pt] e_{2,1}^{a_{2,1}}\,e_{4,3}^{a_{4,3}+1}v\;-\;a_{2,1}\,e_{2,1}^{a_{2,1}-1}\,e_{2,3}\,e_{4,1}\,e_{4,3}^{a_{4,3}}v,&(a_{2,3},a_{4,1})=(0,1),\\[4.0pt] -\,e_{2,1}^{a_{2,1}}\,e_{2,3}\,e_{4,3}^{a_{4,3}+1}v,&(a_{2,3},a_{4,1})=(1,1).\end{cases}$
Moreover, every term on the right-hand side is written in the PBW order $e_{2,1}e_{2,3}e_{4,1}e_{4,3}$.
By the computation, one directly checks that
$\operatorname{ch}\operatorname{DS}_{e_{1,3}}\operatorname{BG}(\lambda)=e^{\lambda}.$
For a $\operatorname{BG}$-module $\operatorname{BG}(M)$, it is easy to determine whether it is annihilated by $\operatorname{DS}_{e_{1,3}}$.
By repeatedly applying Hinich’s lemma, the above results (1)-(8) can be recovered directly.
Example 6.3.
Work in $U(\mathfrak{g})$ with $\mathfrak{g}=\mathfrak{gl}(2|2)$. Let
$\lambda=a\varepsilon_{1}+b\varepsilon_{2}-b\delta_{1}-c\delta_{2}.$
Then the subsuperalgebra $()+(1)$ admits a one-dimensional module $k_{\lambda}$ on which $\mathfrak{h}$ acts by $\lambda$. Moreover, there is a short exact sequence
$0\to\operatorname{Ind}_{()+(1)}^{\mathfrak{g}}k_{\lambda-\varepsilon_{2}+\delta_{1}}\to M^{()}(\lambda)\to\operatorname{Ind}_{()+(1)}^{\mathfrak{g}}k_{\lambda}\to 0.$
The module $\operatorname{Ind}_{()+(1)}^{\mathfrak{g}}k_{\lambda}$ admits a PBW-type basis consisting of vectors
$e_{2,1}^{a_{2,1}}e_{3,1}^{a_{3,1}}e_{4,1}^{a_{4,1}}e_{4,2}^{a_{4,2}}e_{4,3}^{a_{4,3}}v,$
where $v$ is a highest weight vector. Take
$a_{2,1},a_{4,3}\in\mathbb{Z}_{\geq 0},\qquad a_{3,1},a_{4,1},a_{4,2}\in\mathbb{Z}/2\mathbb{Z}.$
Then one computes
$e_{2,3}\,e_{2,1}^{a_{2,1}}e_{3,1}^{a_{3,1}}e_{4,1}^{a_{4,1}}e_{4,2}^{a_{4,2}}e_{4,3}^{a_{4,3}}v$
$\qquad=\begin{cases}0,&(a_{3,1},a_{4,2})=(0,0),\\[4.0pt] e_{2,1}^{a_{2,1}+1}\,e_{4,1}^{a_{4,1}}\,e_{4,3}^{a_{4,3}}v,&(a_{3,1},a_{4,2})=(1,0),\\[4.0pt] (-1)^{a_{4,1}}\,e_{2,1}^{a_{2,1}}\,e_{4,1}^{a_{4,1}}\,e_{4,3}^{a_{4,3}+1}v,&(a_{3,1},a_{4,2})=(0,1),\\[4.0pt] e_{2,1}^{a_{2,1}+1}\,e_{4,1}^{a_{4,1}}\,e_{4,2}\,e_{4,3}^{a_{4,3}}v\\ \qquad\;-\;(-1)^{a_{4,1}}\,e_{2,1}^{a_{2,1}}\,e_{3,1}\,e_{4,1}^{a_{4,1}}\,e_{4,3}^{a_{4,3}+1}v,&(a_{3,1},a_{4,2})=(1,1).\end{cases}$
Moreover, every term on the right-hand side is written in the PBW order $e_{2,1}e_{3,1}e_{4,1}e_{4,2}e_{4,3}$.
It follows that $\operatorname{DS}_{e_{2,3}}\operatorname{Ind}_{()+(1)}^{\mathfrak{g}}k_{\lambda}$ is spanned by $v$ and $e_{4,1}v$. Consequently,
$\operatorname{DS}_{e_{2,3}}\operatorname{Ind}_{()+(1)}^{\mathfrak{g}}k_{\lambda}\cong\Pi^{\operatorname{par}({b|b})}M^{()}(a\mid c).$
By Hinich’s lemma, there is an exact sequence
$0\to\Pi E\to\Pi\,\Pi^{\operatorname{par}({b|b})}M^{()}(a\mid c)\to\operatorname{DS}_{e_{2,3}}M^{()}(\lambda)\to\Pi^{\operatorname{par}({b|b})}M^{()}(a\mid c)\to E\to 0.$
Since the highest weight vector is not killed, we have $\operatorname{DS}_{e_{2,3}}M^{()}(\lambda)\neq 0$. There is no nonzero object $E$ such that both $\operatorname{Hom}(E,M^{()}(a\mid c))\neq 0$ and $\operatorname{Hom}(M^{()}(a\mid c),E)\neq 0$. Therefore $E$ in Hinich’s lemma is zero, hence
$\operatorname{DS}_{e_{2,3}}M^{()}(\lambda)\cong M^{()}(a\mid c)\oplus\Pi\,M^{()}(a\mid c).$
Example 6.4.
Work in $U(\mathfrak{gl}(2|2))$ with $|1|=|2|=\bar{0}$ and $|3|=|4|=\bar{1}$.
Let
$\lambda=a\varepsilon_{1}+b\varepsilon_{2}-b\delta_{1}-c\delta_{2}.$
Then the subsuperalgebra $()+(1)$ admits a one-dimensional module $k_{\lambda}$ on which $\mathfrak{h}$ acts by $\lambda$. Moreover, there is a short exact sequence
$0\to\operatorname{Ind}_{()+(1)}^{\mathfrak{g}}k_{\lambda+\varepsilon_{2}-\delta_{1}}\to M^{(1)}(\lambda)\to\operatorname{Ind}_{()+(1)}^{\mathfrak{g}}k_{\lambda}\to 0.$
The module $\operatorname{Ind}_{()+(1)}^{\mathfrak{g}}k_{\lambda}$ admits a PBW-type basis consisting of vectors
$e_{2,1}^{a_{2,1}}e_{3,1}^{a_{3,1}}e_{4,1}^{a_{4,1}}e_{4,2}^{a_{4,2}}e_{4,3}^{a_{4,3}}v,$
where $v$ is a highest weight vector. Take
$a_{2,1},a_{4,3}\in\mathbb{Z}_{\geq 0},\qquad a_{3,1},a_{4,1},a_{4,2}\in\mathbb{Z}/2\mathbb{Z},$
and assume $e_{3,2}v=0$. Then
$e_{3,2}\,e_{2,1}^{a_{2,1}}e_{3,1}^{a_{3,1}}e_{4,1}^{a_{4,1}}e_{4,2}^{a_{4,2}}e_{4,3}^{a_{4,3}}v$
$\qquad=\begin{cases}0,&(a_{3,1},a_{4,2})=(1,1),\\[4.0pt] a_{2,1}\,e_{2,1}^{a_{2,1}-1}\,e_{3,1}\,e_{4,1}^{a_{4,1}}e_{4,3}^{a_{4,3}}v\;-\;a_{4,3}(-1)^{a_{4,1}}\,e_{2,1}^{a_{2,1}}e_{4,1}^{a_{4,1}}\,e_{4,2}\,e_{4,3}^{a_{4,3}-1}v,&(a_{3,1},a_{4,2})=(0,0),\\[6.0pt] a_{2,1}\,e_{2,1}^{a_{2,1}-1}\,e_{3,1}\,e_{4,1}^{a_{4,1}}\,e_{4,2}\,e_{4,3}^{a_{4,3}}v,&(a_{3,1},a_{4,2})=(0,1),\\[6.0pt] a_{4,3}(-1)^{a_{4,1}}\,e_{2,1}^{a_{2,1}}\,e_{3,1}\,e_{4,1}^{a_{4,1}}\,e_{4,2}\,e_{4,3}^{a_{4,3}-1}v,&(a_{3,1},a_{4,2})=(1,0).\end{cases}$
Moreover, every term on the right-hand side is written in the PBW order $e_{2,1}e_{3,1}e_{4,1}e_{4,2}e_{4,3}$. We can check
$\operatorname{DS}_{e_{3,2}}\operatorname{Ind}_{()+(1)}^{\mathfrak{g}}k_{\lambda}\cong\Pi^{\operatorname{par}({b|b})}M^{()}(a\mid c).$
We also see that $E$ in Hinich’s lemma is zero, hence
$\operatorname{DS}_{e_{3,2}}M^{(1)}(\lambda)\cong M^{()}(a\mid c)\oplus\Pi\,M^{()}(a\mid c).$
Definition 6.5.
Let $\alpha=\varepsilon_{i}-\varepsilon_{j}$ be an odd root with root vector $e_{\alpha}=e_{i,j}$, where $1\leq i\leq n<j\leq 2n$. Define
$\operatorname{pr}_{\alpha}:\bigoplus_{t=1}^{2n}\mathbb{Z}\,\varepsilon_{t}\longrightarrow\bigoplus_{\begin{subarray}{c}1\leq t\leq 2n\\ t\neq i,j\end{subarray}}\mathbb{Z}\,\varepsilon_{t}$
by $\operatorname{pr}_{\alpha}(\varepsilon_{i})=\operatorname{pr}_{\alpha}(\varepsilon_{j})=0$ and $\operatorname{pr}_{\alpha}(\varepsilon_{t})=\varepsilon_{t}$ for $t\neq i,j$.
Combining the results of this section, together with the Lie superalgebra automorphisms $(\cdot)^{c}$ and $(\cdot)^{\operatorname{at}}$, we obtain the following.
Theorem 6.6.
Let $\mathfrak{g}=\mathfrak{gl}(2|2)$, let $\mathfrak{b}\in L(2,2)$, and let $\alpha$ be a $\mathfrak{b}$-simple odd root. Then, for $\lambda\in\Lambda$, we have
$\operatorname{DS}_{e_{\alpha}}\bigl(M^{\mathfrak{b}}(\lambda)\bigr)\ \cong\ \begin{cases}M^{\mathfrak{b}_{e_{\alpha}}}\bigl(\operatorname{pr}_{\alpha}(\lambda)\bigr)\ \oplus\ \Pi\,M^{\mathfrak{b}_{e_{\alpha}}}\bigl(\operatorname{pr}_{\alpha}(\lambda)\bigr),&\text{if }(\lambda,\alpha)=0,\\[5.69054pt] 0,&\text{if }(\lambda,\alpha)\neq 0.\end{cases}$